OpenStudy (anonymous):

find the slope and y intercept of each line of the given equation -4x+y = 7

OpenStudy (anonymous):

We need to get this into slope-intercept form, y=mx+b, where m=slope and b=y-intercept. -4x +y = 7 We'll keep y on the left and move -4x to the right. -4x +4x +y = 7 +4x y = 4x +7 This is now in form y=mx+b. You can pull out the m and b values for your answer.

OpenStudy (anonymous):

so 4 and 7?